OpenStudy (anonymous):

Find all solutions to the equation. cos^2x + 2 cos x + 1 = 0

OpenStudy (anonymous):

(cos(x) +1)(cos(x)+1) = 0 which is true whenever cosx =-1

OpenStudy (anonymous):

@Algebraic! can you explain how you did that?

OpenStudy (anonymous):

factored cos^2x + 2 cos x + 1

OpenStudy (anonymous):

expand (cos(x) +1)(cos(x)+1) and you'll see it's equivalent.

OpenStudy (anonymous):

or... pretend you're factoring u^2 +2u +1 (where u = cos(x) )

OpenStudy (anonymous):

@Algebraic! Thank you

OpenStudy (anonymous):

sure!